OpenStudy (anonymous):

Two apples and three pears cost $3.45. Three apples and five pears cost $5.55. Find the cost of each type of fruit.

OpenStudy (anonymous):

well you have the equation 2a + 3p = 3.45 3a + 5p = 5.55 so if you have 1 apple and 1 pear how much would you have to pay??

OpenStudy (anonymous):

There are your equations. What you want to do is find a way to eliminate a variable. 2a + 3p = 3.45 3 * (2a + 3p) = 3 * 3.45 6a + 9p = 10.35 3a + 5p = 5.55 2 * (3a + 5p) = 2 * 5.55 6a + 10p = 11.10

OpenStudy (anonymous):

so now because we have 6a in both equations we must now find out the answer. (6a + 10p) - (6a + 9p) = 11.10 - 10.35 6a - 6a + 10p - 9p = 0.75 p = 0.75 So, 1 pear costs $0.75 2a + 3p = 3.45 2a + 3 * 0.75 = 3.45 2a + 2.25 = 3.45 2a = 1.2 a = 0.6 1 apple costs $0.60

OpenStudy (anonymous):

Thank you